OpenStudy (anonymous):

five time the sum of a number and -3

OpenStudy (anonymous):

??? please give the complete question

OpenStudy (anonymous):

use a varible x to represent the unknown number in each expression unless otherwise noted You have an open question: five time the sum of a number and -3Asked 2 minutes ago 1 reply No Medals Yet Bump in 2:37Close Open Questions Closed Questions 29LovePink:Reagan has a savings account that earns 1.87% interest, compounding quarterly. If she needs $20,857 for a down payment on a house in nine ye… 1 viewing No Medals Yet 17immeen04:i am having a difficult time solving this differential equation: (dy/dx) = sin (x + y) + cos(x +y) 1 viewing No Medals Yet 52viodhora:Solve equation (2+2i) z^3=-1, where z is complex number. 2 viewing No Medals Yet 15FishingWithTNT:Twice one number added to another is 100. What should the numbers be if the sum of their squares is to be a minimum? How do I do this? updated a minute ago No Medals Yet 29Barrington:I Really need help i have to know this by tomorrow when i take my exam i have a hole study guide thats really complicated 1 viewing No Medals Yet 46monroe17:Let U={q,r,s,t,u,v,w,x,y,z}, A={q,s,u,w,y}, and B={q,s,y,z}. List the elements in the set A′∪B. …1 replying right now... No Medals Yet 8extreme34810:five time the sum of a number and -3 1 viewing No Medals Yet 61Answerthisnow:Triangle XYZ is shown on the coordinate plane below. … 1 viewing No Medals Yet 23suzan998:What is the smallest multiple of 9997 (other than 9997 itself) which contains only odd digits? 4 viewing No Medals Yet 1109rainbowc:5(2x-7) =3 updated 6 minutes ago 1 Medal No more questions. Working hard or hardly working? Privacy Policy Terms and Conditions Code of Conduct annas says... Improve Your Answer You can add equations, attachments, and drawings to better explain your answers. close Don't show tipsClose Search Subjects Mathematics 56123 members, 334551 questions 72 Online Stop Studying Physics 8688 members, 13764 questions 38 Online Writing 6993 members, 7069 questions 27 Online OpenStudy Feedback 2534 members, 2831 questions 23 Online Computer Science 6254 members, 4678 questions 22 Online Chemistry 3818 members, 6732 questions 21 Online Biology 6922 members, 11978 questions 21 Online History 4586 members, 6279 questions 18 Online Music 558 members, 458 questions 14 Online Engineering 1077 members, 427 questions 10 Online Loading more… Create a New Subject Enter Subject Name Cancel

OpenStudy (anonymous):

sounds like an exponent SITUATION ! WHATS THE QUESTION ?

OpenStudy (anonymous):

let a number is x 5 times of x and -3 is: 5(x-3)